OpenStudy (anonymous):

( -5) TWO SECOND POWER

OpenStudy (anonymous):

please show work been out of school for 20 years . jus need a step by step example

OpenStudy (amistre64):

a power tells us how many times we are going to multiply this number by itself to get our answer. for example: x^5 means x*x*x*x*x = x^5 (-5)^2 means (-5)(-5) = 25 If you had written -(5)^2 it would be asking for the opposite of (5)^2. which would be the opposite if (5)(5) = -25

OpenStudy (anonymous):

thanks so much ; i really do appericated your

OpenStudy (anonymous):

help . now i feel like i can go on . thanks so much

OpenStudy (amistre64):

trust me, if you know how to count to (1) the rest will come naturally :) Math is an abstract art and so it needs to develop a language to describe the simple things we do in life and make then very complex. But in the end, all you are doing is counting :)